OpenStudy (anonymous):

please help. I need to sketch a graph of y=f(x) f(x)=-4x+2

terenzreignz (terenzreignz):

Graphing lines? :)

OpenStudy (anonymous):

yes please

terenzreignz (terenzreignz):

...

terenzreignz (terenzreignz):

Graphing lines almost always involve finding two points on the line and just connecting them :) Typically, the two easiest points to find are the intercepts. Do you know what intercepts are? ^_^

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well then, what is the x-intercept and y-intercept of this line?

OpenStudy (anonymous):

sorry I'm still trying to solve f(x)=-4+2

terenzreignz (terenzreignz):

You mean f(x) = -4x + 2 And yes, that's what I mean, find the intercepts of that line...

OpenStudy (anonymous):

-4 on the x line

terenzreignz (terenzreignz):

not quite... Tell you what, set x = 0, and whatever value y-gets, that's the y-intercept

OpenStudy (anonymous):

so i take -4(0)+2

terenzreignz (terenzreignz):

Yup, and the value is...?

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

And that... is the y-intercept :) So one of the points on the line is (0,2) Let's mark that out...

terenzreignz (terenzreignz):

Now, find the x-intercept. This time, set f(x) = 0, and then solve for x.

OpenStudy (anonymous):

-4x+2(0)

terenzreignz (terenzreignz):

No, uhh f(x) = -4x + 2 now set f(x) = 0 0 = -4x + 2 and solve for x

OpenStudy (anonymous):

x=2

terenzreignz (terenzreignz):

No... try again

OpenStudy (anonymous):

trying -4x+2=0 x=4

terenzreignz (terenzreignz):

-4x = -2 now divide both sides by -4

OpenStudy (anonymous):

2

OpenStudy (anonymous):

no-2

OpenStudy (anonymous):

i do mean 2

terenzreignz (terenzreignz):

It's neither... come on, if x = 2, then -4(2) = -2 -8 = -2 not true

OpenStudy (anonymous):

so i was right with -2...I am horrible with my - signs

terenzreignz (terenzreignz):

-4(-2) = 8 which is still not -2

OpenStudy (anonymous):

sorry i hate algebra

OpenStudy (anonymous):

so x=8 and my lines are (8,0) for x and (0,2) for y

terenzreignz (terenzreignz):

No... the x intercept would be whatever value you get for x in -4x = -2 which you still haven't solved yet :P

OpenStudy (anonymous):

-2/1

terenzreignz (terenzreignz):

That's the same as -2?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

wrong

OpenStudy (anonymous):

sorry i can't do this i give up thank you anyways